1. The Hardaway Corporation plans to lease a $770,000 asset to the O’Neil Corporation. The lease...

1.

The Hardaway Corporation plans to lease a $770,000 asset to the O’Neil Corporation. The lease will be for 16 years. Use Appendix D for an approximate answer but calculate your final answer using the formula and financial calculator methods.

a. If the Hardaway Corporation desires a return of 13 percent on its investment, how much should the lease payments be? (Do not round intermediate calculations and round your answer to 2 decimal places.)
  

Lease payment

b. If the Hardaway Corporation is able to take a 10 percent deduction from the purchase price of $770,000 and will pass the benefits along to the O’Neil Corporation in the form of lower lease payments (related to the Hardaway Corporation in the form of lower initial net cost), how much should the revised lease payments be? The Hardaway Corporation desires a return of 13 percent on the 16-year lease. (Do not round intermediate calculations and round your answer to 2 decimal places.)
  

Revised lease payment

Q2.

A previously issued A2, 15-year industrial bond provides a return three-fourths higher than the prime interest rate of 14 percent. Previously issued A2 public utility bonds provide a yield of five-eighths of a percentage point higher than previously issued A2 industrial bonds of equal quality. Finally, new issues of A2 public utility bonds pay three-fourths of a percentage point more than previously issued A2 public utility bonds.


What should be the interest rate on a newly issued A2 public utility bond? (Do not round intermediate calculations. Input your answer as a percent rounded to 3 decimal places.)
  

Interest rate

%

Q3.

The Bowman Corporation has a bond obligation of $18 million outstanding, which it is considering refunding. Though the bonds were initially issued at 11 percent, the interest rates on similar issues have declined to 9.8 percent. The bonds were originally issued for 20 years and have 10 years remaining. The new issue would be for 10 years. There is a call premium of 7 percent on the old issue. The underwriting cost on the new $18,000,000 issue is $480,000, and the underwriting cost on the old issue was $370,000. The company is in a 35 percent tax bracket, and it will use an 8 percent discount rate (rounded aftertax cost of debt) to analyze the refunding decision. Use Appendix D for an approximate answer but calculate your final answer using the formula and financial calculator methods.

a. Calculate the present value of total outflows.(Do not round intermediate calculations and round your answer to 2 decimal places.)
  

PV of total outflows

b. Calculate the present value of total inflows.(Do not round intermediate calculations and round your answer to 2 decimal places.)
  

PV of total outflows

c. Calculate the net present value. (Negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places.)
  

Net present value
Answer #1
1. a.The annual Lease payment can be found out by using the Present value of annuity formula,
PV of the lease=Annuity amt.*(1-(1+r)^-n)/r
where, PV of the lease is known,ie. $ 770000
Annuity amt.-- needs to be found out --??
r=Rate of interest per annum,ie. 13% p.a.
n=no.of compounding periods= 16
So, plugging in the known values.
770000=Annuity Amt.*(1-1.13^-16)/0.13
& solving for the amt., we get the annuity amt. as
770000/((1-1.13^-16)/0.13)=
116598.21
ANSWER:
The annual lease payment = $ 116598.21
b.Using the same formula as in a. above,
the revised lease pmt. Will be
(770000*(1-10%))=Annuity Amt.*(1-1.13^-16)/0.13
104938.39
ANSWER:
The revised annual lease payment = $ 104938.39
2..Return on A2 ,15 yr, industrial bond(14%+(14%*3/4)= 24.5%
Plus:Current Return on previously issued A2 P/U bond(1%*5/8) 0.625%
Plus:Return on new issues of A2 P/U bond(1%*3/4) 0.75%
Interest rate on a newly issued A2 public utility bond 25.875%
Cash outflows to refund the issue :
After-tax Call premium on old issue
18000000*7%*(1-35%)= -819000
Underwriting costs on costs on new issue -480000
Unamortised Underwriting costs on costs on old issue
(370000/20*10*35%) 64750
Net outflows -1234250
Present value of Inflows: P/A,8%,10 PV
Tax savings on U/w costs of the new issue
480000/10*35% 16800 6.71008 112729
Tax benefits lost on U/w costs of old issue
185000/10*35% -6475 6.71008 -43448
After-tax interest savings on annual coupons
18000000*(11%-9.8%)*(1-35%) 140400 6.71008 942095
PV of net inflows 1011377
P/A,i=8%, n=10 yrs.
6.71008
NPV of the refunding -222873